OpenStudy (anonymous):

Calculate the volume of a prism with measurements a = 11, b = 5, and c = 10. 550 450 325 275

OpenStudy (anonymous):

Volume = abc (assuming that a,b, and c are the lengths of the prism's sides) V = (11)(5)(10) = 550

OpenStudy (anonymous):

Here is the prism.

OpenStudy (anonymous):

i took a quiz with the same question on it and the answer was 275. I thought that was wrong but idk.

OpenStudy (anonymous):

i keep on getting 550

OpenStudy (anonymous):

Ok to find the volume, we find the area of the triangle and then multiply it by b.

OpenStudy (anonymous):

Area of a triangle = (1/2)bh = (1/2)(10)(11) = 55

OpenStudy (anonymous):

Multiply by 5 to get 275

OpenStudy (anonymous):

Notice that 275 is half of 550, we just didn't note that it's a triangle and not a rectangular prism.

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

thank you, i didnt understand it before

OpenStudy (anonymous):

happy to help :)

OpenStudy (anonymous):

can you help me with another one? its harder and no one can help. I'm so frustrated with it.

OpenStudy (anonymous):

I can try